4. In an experiment, 9.80 g of olive oil at 80.0C are quickly transferred to water in a calorimeter at 18.0°C. The heat capacity of the water is 4.18 J/gC. The final equilibrium temperature is 22.8°C. What is the approximate specific heat capacity of the olive oil?
5. ) Given the standard enthalpies of formation for the following substances, determine the change in enthalpy for the combustion of 1.0 mol of propane, C3H8.
Substance ΔH°f in kJ/mol
C3H8(g) 103.8
CO2(g) 393.5
H2O(l) 285.8
6. Calculate the energy released as heat when 45.8 g of Fe(s) is combined with 27.9g of O2(g) if the ΔH°rxn is -1648 kJ/mol. The balanced equation for the reaction is
4 Fe(s) + 3O2(g) = 2 Fe2O3(s)
1
Expert's answer
2019-12-12T07:30:50-0500

4. Q=mc∆T = 9.8 x 4.18 x 4.8 = 196.6272 J.

196.6272/9.8= 20.064 J/gC.

5.

C3H8 + 5O2 = 3CO2 + 4H2O.

4 x 285.8 + 3 x 393.5 - 103.8 = 2,219.9 kJ/mol.

6.

4 Fe(s) + 3O2(g) = 2 Fe2O3(s)

96 g --------- - 1648 .

27.9 --------- x

x= 27.9 x 1648/96= - 478.95 kJ/mol.